Jamie Lee knows how hard it is to go to school, take on debt, and try to fulfill her dreams. Her sister, Rebecca, recently had a baby girl named Alexis. Jamie Lee would like to set aside a small amount per month towards Alexiss education expenses. Jamie Lee realizes this will not cover all of the costs but, she wants to help in a small way After researching college costs and investments, Jamie Lee has the following information: Current cost of college education: $32,500 per year (Alexis will attend college in 18 years for 4 years.) Expected annual inflation: 2% Expected rate of return on investment: 11% Use the table below to determine the estimated annual deposit needed to achieve the education fund. (Use Exhibit 1-A Exhibit 1-8) Estimated Cost of College Education Current cost of college education (including tuition, fees, room, board, books, travel and other expenses) Years until starting college Expected annual inflation Future value factor (use Exhibit 1-A in Chapter 1 Appendix) Projected future cost of college adjusted for inflation 0.00 Estimated Annual Savings Needed Projected future cost of college adjusted for inflation Years until starting college Expected annual rate of return on savings and investments Future value of a series of deposits factor (use Exhibit 1-B in Chapter 1 Appendix) Estimated annual deposit to achieve needed educational fund 0.00 NaN
Answer #1
Estimated Cost of College Education:
A Current Cost of college education per year $            130,000 (32500*4)
B Years untill   starting college 18
C Expected annual inflation 2%
D=(1+C)^B Future Value Factor 1.428246248 (1+0.02)^18
E=A*D Projected future cost(inflation adjusted) $            185,672
Estimated Annual Saving Needed:
E Projected future cost(inflation adjusted) $            185,672
B Years until starting college 18
F Expected annual return on savings/investment 11%
G Future Value of series of deposit factor           0.0198429 (Using PMT function of excel with Rate=11%,Nper=18, Fv=-1)
H=E*G Estimated Annual Deposit needed $3,684